DETECTION OF VAN DER WAALS VIBRATIONAL TRANSITIONS USING A NEW MW-THz DOUBLE RESONANCE TECHNIQUE

VLADIMIR MARKOV, ^{a} YUNJIE XU, and \underline{WOLFGANG JÄGER}, Department of Chemistry, University of Alberta, Edmonton, AB, Canada, T6G 2G2.

A new double resonance technique has been developed that can be used for the detection of low frequency van der Waals vibrations. The technique uses backward wave oscillators (OB-24, OB-30, ISTOK), which are phase-locked to a KVARZ mm-wave synthesizer (78-118 GHz), as pump sources, ^{b} and a pulsed jet cavity Fourier transform microwave spectrometer for detection. Initial measurements were done on the Ar-CO system ^{c} for which the THz frequencies of ro-vibrational transitions from the ground to the first excited van der Waals bending state were known from previous measurements. ^{d} Resolving power and sensitivity of the technique will be discussed.

apermanent address: Institute of Applied Physics of the Russian Academy of Sciences, Ulyanova St. 46, 603600 Nizhnii Novgorod, Russia.

bG. Winnewisser, A. F. Krupnov, M. Yu. Tretyakov, M. Liedtke, F. Lewen, A. H. Salek, R. Schieder, A. P. Shkaev, and S. V. Volokhov, J. Mol. Spectrosc. 165, 294 (1994).

cY. Xu and A. R. W. McKellar, Mol. Phys. 88, 859 (1996).

dM. Hepp, R. Gendriesch, I. Pak, F. Lewen, and G. Winnewisser, J. Mol. Spectrosc. 183, 295 (1997).